OpenStudy (anonymous):

Find the zeros of the function f(x) = x^3 + 6x^2 + 15x + 10

OpenStudy (anonymous):

Hello?

OpenStudy (wolf1728):

Wow you like answers fast!!

OpenStudy (anonymous):

hahahaha sorry

OpenStudy (wolf1728):

Here they are: -1 -2.5 + i* 1.9364916731037085 -2.5 - i* 1.9364916731037085

OpenStudy (wolf1728):

I'd explain how I got them but I don't want to waste any more time LOL !!!!!!!

Parth (parthkohli):

Observing the function, \(x^3 + 6x^2 + 15x\) has to be \(-10\) so as to make the whole expression zero. From observation, you can see how \(-1 + 6 - 15 = -10\), thus making \(-1\) a root of the expression.

mathslover (mathslover):

Yes, right! First find a solution ........

OpenStudy (anonymous):

is -1 the only solution?

mathslover (mathslover):

Let you have three solutions for this : \(\alpha , \beta ,\gamma \) then : \((x-\alpha ) (x- \beta)( x- \gamma) = 0 \) We got \(\alpha = -1\) Thus, \((x+1)(x -\beta)(x - \gamma) = 0 \) Clearly : \((x-\beta)(x-\gamma) \) can be written as : \(x^2 + ax + b\) where a and b are numbers.

OpenStudy (anonymous):

Thanks you all!

mathslover (mathslover):

Use synthetic division now, to find a and b : Write down the coefficients of the original expression as : 1 6 15 10 | x = -1 | ----------------------------------------- 1 I have written down the first coefficient (1) Now multiply the "1" by "-1" = -1 Write this "-1" below the number "6" 1 6 15 10 | x = -1 -1 ----------------------------------------- 1 Add 6 and -1 1 6 15 10 | x = -1 -1 ----------------------------------------- 1 5 Now, again, multiply the "5" with "-1" : 1 6 15 10 | x = -1 -1 -5 ----------------------------------------- 1 5 Add 15 and -5 1 6 15 10 | x = -1 -1 -5 ----------------------------------------- 1 5 10 Similarly, multiply 10 with -1 1 6 15 10 | x = -1 -1 -5 -10 ----------------------------------------- 1 5 10 Add 10 and -10 1 6 15 10 | x = -1 -1 -5 -10 ----------------------------------------- 1 5 10 0 As the last answer we got is 0, thus "x = -1" is one root of the given equation. Now, note that those : \(\bf{ 1 \space , 5 \space , 10 \space }\) are the coefficients of the expression : \(x^2 + ax + b\) Thus, you get : \(x^2 + ax + b = 1(x^2) + 5(x) + 10\) Above, we had : \((x+1)(x^2 + ax+ b) = 0 \) Thus, it becomes : \((x+1)(x^2 + 5x + 10) =0 \) One solution is x + 1 = 0 or x = - 1 Second and third solution will be obtained from : \(x^2 + 5x + 10 =0 \) \(x = \cfrac{-5 \pm \sqrt{25 - 40}}{2} = \cfrac{ -5 \pm \sqrt{15} i }{2} \) Or : x = \(\cfrac{-1}{2} \left( -5 + i\sqrt{15} \right) \) or x = \(\cfrac{-1}{2} \left( -5 - i\sqrt{15} \right) \) or x = -1. I hope it helps.